OpenStudy (anonymous):

given y=4x^2-2x+(2), find the equation of the tangent line that is horizontal to the curve

sam (.sam.):

You can differentiate it first

OpenStudy (anonymous):

and you get 8x-2

OpenStudy (anonymous):

what do you do after

sam (.sam.):

you can try equating dy/dx=0 then solve for x

sam (.sam.):

dy/dx=8x-2 8x-2=0 x=1/4

sam (.sam.):

so y=4x^2-2x+(2) y=4(1/4)^2-2(1/4)+2 y=7/4

OpenStudy (anonymous):

how do we put it in slope intercept form?

sam (.sam.):

then use \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

whats m=?

OpenStudy (anonymous):

so the final answer is?

sam (.sam.):

equation of tangent line is y=7/4

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can you help me with one more problem?